if the diameter of a newly discovered atom is 3.94 × 10-8cm, how many of these atoms would fit along a 2.78in line
Answer #1

Diameter of an atom = 3.94*10^-8 cm

Length of line = 2.78 in

= 2.78 * 2.54 cm [Since 1 inch = 2.54 cm]

= 7.06 cm

Use:

Number of atoms = Length of line / Diameter of an atom

= 7.06 / (3.94*10^-8)

= 1.79*10^8

Answer: 1.79*10^8